# Sampling as Bandits: Evaluation-Efficient Design for Black-Box Densities

### v1

@@ 3 Methodology
Suppose $\Theta\subset\mathbb{R}^{d}$ is a compact hyperrectangular domain with no loss of generality, since reparametrisation can be performed otherwise. Let $p(\theta)=q(\theta)/Z$ be a target density of interest on $\Theta$, where $q(\theta)$ denotes the proportional term and $Z$ denotes the normalisation constant. Assume that $p(\theta)$ is positive over $\Theta$. A practical scenario in mind is that evaluation of the proportional $q(\theta)$ is possible at each $\theta$ but computationally expensive for one’s computational capacity.
@@ 3.1 Bandit Importance Sampling
The aim of BIS is to produce a set of $N$ weighted samples, $\{\theta^{*}_{n},w^{*}_{n}\}_{n=1}^{N}$, which approximates the target density $p(\theta)$, requiring only $N$ evaluations of the proportional $q(\theta)$ in total. Rather than drawing samples randomly from a proposal distribution, BIS directly optimises the samples themselves. BIS selects each sample sequentially from a pool of candidates under a chosen criterion. This can be viewed as a multi-armed bandit problem, where at each iteration a new sample is selected based on current information. Notably, once a point is selected in BIS, it is not revisited in future iterations (see Remark˜1), distinguishing it from conventional bandit formulations.
We detail the iterative procedure of BIS. Let $u(\theta)$ be a proposal density on $\Theta$. Let $\mathcal{U}=\{\theta_{i}\}_{i=1}^{\infty}$ be a sequence that weakly converges to $u(\theta)$, which we call a proposal sequence. Let $S_{n}$ denote a pool of candidates at each iteration $n$. At the first iteration $n=1$, we use the first $M$ points of the proposal sequence $\mathcal{U}$ as the selection pool $S_{1}=\{\theta_{i}\}_{i=1}^{M}$. The candidate pool $S_{n}$ is updated after each iteration, replacing the selected point with a new point from the proposal sequence. This assures no point in the proposal sequence $\mathcal{U}$ will be selected twice. Finally, let $U_{n}(\theta)$ be a point-selection criterion specified at each iteration $n$. We repeat the following steps from $n=1$:
@@
    choose a point $\theta_{n}^{*}$ in the candidate pool $S_{n}$, which maximises the criterion $U_{n}(\theta)$;
@@
    evaluate the target proportional $q(\theta_{n}^{*})$ for evaluation of the importance weight $w_{n}^{*}$;
@@
    specify the candidate pool $S_{n+1}$ at the next iteration by $S_{n+1}=(S_{n}\setminus\{\theta_{n}^{*}\})\cup\{\theta_{M+n}\}$.
We assign the self-normalised weight $w_{n}^{*}=(q(\theta_{n}^{*})/u(\theta_{n}^{*}))/\sum_{i=1}^{N}(q(\theta_{i}^{*})/u(\theta_{i}^{*}))$ to each of the samples $\{\theta_{n}^{*}\}_{n=1}^{N}$ after the last iteration $N$. Note that the size of the candidate pool, $|S_{n}|$, is constant at $|S_{n}|=M$ across all iterations. Algorithm 1 summarises the explicit procedure.
@@
$S_{1}\leftarrow\{\theta_{i}\}_{i=1}^{M}$ $\triangleright$ initialise selection pool
@@
 $\theta_{n}^{*}\leftarrow\operatorname*{arg\,max}_{\theta\in S_{n}}U_{n}(\theta)$ $\triangleright$ optimise point-selection criterion
@@
$(w_{1}^{*},\cdots,w_{N}^{*})\leftarrow(w_{1}^{*},\cdots,w_{N}^{*})/\sum_{n=1}^{N}w_{n}^{*}$ $\triangleright$ normalise weights
@@
Remark 1 (Importance of Discrete Optimisation with No Revisit).
BIS selects a point by optimisation of the criterion $U_{n}$ over the discrete set $S_{n}$, under the constraint that points selected in the past will never be revisited. A possible alternative, similarly to BO, is to select a point by optimisation of the criterion $U_{n}$ over the continuous set $\Theta$. However, as we shall demonstrate in Section˜C.1, the latter approach can situate an excessive portion of samples around the mode of the target density. Consequently, this derails the convergence of the resulting weighted samples, in addition to causing redundant density evaluations around the mode. On the other hand, the use of the discrete set $S_{n}$ in BIS serves as a regularisation to prevent overconcentration of the selected points. In combination with that, the mechanism of not revisiting the past points ensures the convergence of the samples to the target posterior (c.f. Section˜4).
@@ 3.2 Proposal Density and Proposal Sequence
We discuss a choice of the proposal density $u$ and the proposal sequence $\mathcal{U}=\{\theta_{i}\}_{i=1}^{\infty}$ in BIS. In standard importance sampling, there exists an optimal proposal density in terms with the variance reduction of the Monte Carlo estimation. However, such a proposal density depends on the target density itself, which is black-box and not available in our setting. In the first place, the lack of information on the target density—such as the mode and the decay of the tail probability—means that we have little knowledge with which to design the proposal density. We therefore consider the uniform density over the domain $\Theta$ as a natural default choice of the proposal density $u$.
Note that the proposal sequence $\mathcal{U}$ does not have to be i.i.d. random samples from the proposal density $u$, since unbiasedness is no longer of interest in BIS. Any dependent random sequence or deterministic sequence can be used if it converges to $u$. In this regard, space-filling sequences in QMC exhibit fast convergence to the uniform proposal density $u$ (Morokoff and Caflisch, 1994). Our default choice of the proposal sequence is the Halton sequence (Halton, 1964), which is arguably the most common space-filling sequence in practice. It is originally defined for the unit-cube domain but is immediately amenable to hyperrectangular domains. For clarity, we define a version of the Halton sequence scaled to the hyperrectangular domain $\Theta=[a_{(1)},b_{(1)}]\times\cdots\times[a_{(d)},b_{(d)}]$, where $a_{(i)}$ and $b_{(i)}$ denote the left and right boundary of the domain at the $i$-th coordinate.
@@
Given the Halton sequence $\{\eta_{i}\}_{i=1}^{\infty}$ on the unit cube $0,1]^{d}$, the scaled Halton sequence $\{\theta_{i}\}_{i=1}^{\infty}$ on the domain $\Theta$ is defined by $\theta_{i}:=A+B\eta_{i}$, where $A:=(a_{(1)},\cdots,a_{(d)})\in\mathbb{R}^{d}$ and $B=\mbox{diag}(b_{(1)}-a_{(1)},\dots,b_{(d)}-a_{(d)})$ is a $d\times d$ diagonal matrix.
Using the Halton sequence in BIS is an effective approach to approximation of black-box densities. Regardless of sampling methods, tracking down the complete geometry of the black-box density necessitates the density to be evaluated everywhere in the domain $\Theta$. The Halton sequences provide a sophisticated design of points that explores the entire domain. BIS then selects points around the high-probability region of the target density at the early iterations by the bandit strategy.
@@ 3.3 Point-Selection Criterion
The point-selection criterion $U_{n}$ focused in this paper is built upon a GP surrogate model of the target density $p(\theta)=q(\theta)/Z$. Whereas densities take only positive values, GPs take both negative and positive values. A common strategy to fit a GP to the density is to apply the log-transformation to the value of the density. More generally, any other transformations that map positive values to $\mathbb{R}$ can be used. This can be stated more formally as below.
Let $\phi:\mathbb{R}\to[0,\infty)$ be a convex non-negative function s.t. the inverse $\phi^{-1}$ exists at least for the positive range $(0,\infty)$. Then the proportional $q(\theta)$ of the target density can be expressed as $q(\theta)=\phi(\phi^{-1}(q(\theta)))$. We shall model the transformed term $\phi^{-1}(q(\theta))$ by a GP. For example, choosing the exponential map $\phi(\cdot)=\exp(\cdot)$ recovers the log-transformation $\phi^{-1}(\cdot)=\log(\cdot)$ as the inverse. Now, we introduce our point-selection criterion, GP-UJB, which is equipped with several appealing properties for multi-armed bandits.
@@
$U_{n}(\theta):=\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[\phi\big{(}f(\theta)\big{)}\right]$ (5)
where $f(\theta)$ denotes the value of the GP sample path $f\sim\mathcal{GP}(m_{n},k_{n})$ evaluated at $\theta$.
Driving both exploitation and exploration is a pivotal factor in most bandit problems. GP-UJB gives an upper bound of the ‘least-error’ plug-in approximation $\phi(m_{n}(\theta))$ of the proportional $q(\theta)$ based on the GP posterior mean $m_{n}$. That is, we have $U_{n}(\theta)=\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}[\phi(f(\theta))]\geq\phi(m_{n}(\theta))$ as an immediate consequence of Jensen’s inequality. Therefore, GP-UJB admits the following decomposition into two terms that drive, respectively, exploitation and exploration:
@@
The first term corresponds to the least-error approximation of the proportional $q(\theta)$ achievable by the GP, hence driving exploitation of past information. The second term, known as the Jensen’s gap, is always non-negative due to Jensen’s inequality. It corresponds to the excess deviation between GP-UJB and the least-error GP approximation of the proportional. The second term amplifies according to the level of uncertainty in the GP posterior at each location $\theta$, hence driving exploration of where the GP posterior is uncertain. The second term equals zero at the training points $\{\theta_{i}\}_{i=1}^{n}$, where the GP posterior has no uncertainty.
We can guarantee that the L2 error between GP-UJB and the target proportional $q(\theta)$ is controlled by the average-case interpolation error of the GP regression. Moreover, if an additional regularity condition is met, the former is specifically smaller than the later. This also supports the use of GP-UJB as it converges faster than the concentration of the GP posterior itself. In the following result, $\|\cdot\|_{L_{2}(\Theta)}$ denotes the L2 norm of functions on $\Theta$ under the Lebesgue measure.

### v2

@@ 3 General Framework
Suppose $\Theta\subset\mathbb{R}^{d}$ is a compact hyperrectangular domain with no loss of generality, since reparametrization can be performed otherwise. Let $p(\theta)=q(\theta)/Z$ be a target density of interest on $\Theta$, where $q(\theta)$ denotes the proportional term and $Z$ denotes the normalization constant. Assume that $p(\theta)$ is positive over $\Theta$. A practical scenario in mind is that evaluation of the unnormalized density $q(\theta)$ is possible at each $\theta$ but computationally expensive for one’s computational capacity.
@@ 3.1 Bandit Importance Sampling
BIS aims to construct a set of $N$ weighted samples $\{\theta^{*}_{n},w^{*}_{n}\}_{n=1}^{N}$ to approximate the target density $p(\theta)$, subject to a budget of $N$ evaluations of the unnormalized target $q(\theta)$. Instead of drawing samples randomly from a proposal density, BIS directly optimizes the samples, sequentially selecting them from a candidate pool based on a given selection criterion. Distinct from conventional bandits, once a candidate is selected, it is never revisited in subsequent iterations (see Remark˜1).
We detail the explicit procedure of BIS. Let $u(\theta)$ be a proposal density on $\Theta$. A sequence $\mathcal{U}=\{\theta_{i}\}_{i=1}^{\infty}$, whose empirical distribution converges weakly to $u(\theta)$, is called the proposal sequence in this work. Let $S_{n}$ denote the candidate pool at iteration $n$. At the initial iteration $n=1$, the candidate pool $S_{1}$ is initialized with the first $M$ points of the proposal sequence $\mathcal{U}$: $S_{1}=\{\theta_{i}\}_{i=1}^{M}$. After each iteration, the candidate pool $S_{n}$ is updated by replacing the selected point with the next available point from the proposal sequence $\mathcal{U}$. This mechanism guarantees that no point in $\mathcal{U}$ is ever selected more than once. Finally, let $U_{n}(\theta)$ be the selection criterion defined at iteration $n$, which governs the choice of the sample. We repeat the following procedure from $n=1$:
@@
    Select a point $\theta_{n}^{*}$ from the candidate pool $S_{n}$ that maximizes the criterion $U_{n}(\theta)$;
@@
    Evaluate the unnormalized target $q(\theta_{n}^{*})$ to compute the corresponding weight $w_{n}^{*}$;
@@
    Update the candidate pool for the next iteration by $S_{n+1}=(S_{n}\setminus\{\theta_{n}^{*}\})\cup\{\theta_{M+n}\}$.
Following the final iteration $N$, we assign the self-normalized weight $w_{n}^{*}$ in (2) to each sample $\theta_{n}^{*}$. By construction, the size of the candidate pool $|S_{n}|$ remains constant at $|S_{n}|=M$ throughout the entire procedure. Algorithm 1 provides a formal summary of the BIS procedure.
@@
$S_{1}\leftarrow\{\theta_{i}\}_{i=1}^{M}$ $\triangleright$ initialize selection pool
@@
 $\theta_{n}^{*}\leftarrow\operatorname*{arg\,max}_{\theta\in S_{n}}U_{n}(\theta)$ $\triangleright$ optimize point-selection criterion
@@
$(w_{1}^{*},\cdots,w_{N}^{*})\leftarrow(w_{1}^{*},\cdots,w_{N}^{*})/\sum_{n=1}^{N}w_{n}^{*}$ $\triangleright$ normalize weights
@@
Remark 1 (Importance of Discrete optimization with No Revisit).
BIS selects points by maximizing the criterion $U_{n}$ over the discrete set $S_{n}$, subject to the constraint that previously selected points are never revisited. A possible alternative, analogous to BO, would be to optimize the criterion $U_{n}$ over the continuous set $\Theta$. However, as demonstrated in Section˜C.1, such a continuous approach can lead to an overconcentration of the samples around the modes of the target density. This clustering behavior results in redundant evaluations and hinders the convergence of the importance-sampling approximation. In contrast, the use of the discrete set $S_{n}$ in BIS acts as a regularizer to prevent the overconcentration of the samples. Combined with the no-revisit mechanism, this ensures the convergence of the importance-sampling approximation, c.f., Section˜4.
@@ 3.2 Proposal Density and Proposal Sequence
We discuss the choice of the proposal density $u$ and the corresponding proposal sequence $\mathcal{U}=\{\theta_{i}\}_{i=1}^{\infty}$ in BIS. In standard importance sampling, an optimal proposal density exists in terms with minimizing the variance of the Monte Carlo estimator. However, this proposal density depends on the target density itself, which is black-box and unavailable in our setup. Since key properties of the black-box target density—such as its modes or tail behavior—are unknown a priori, there is even little basis for constructing a tailored proposal density. Therefore, we adopt the uniform density over the domain $\Theta$ as a natural and robust default choice for the proposal density $u$.
Crucially, the proposal sequence $\mathcal{U}$ need not consist of i.i.d. random samples, since unbiasedness is no longer a primary interest for BIS. Any dependent random or deterministic sequence can be employed, provided it converges weakly to the proposal density $u$. Space-filling sequences from Quasi-Monte Carlo (QMC; Dick2010) are ideal for the proposal sequence, exhibiting rapid convergence to the uniform density $u$ (Morokoff1994).
A celebrated result in QMC theory establishes that numerical integration using $N$ points from a space-filling sequence achieves a convergence rate of $\mathcal{O}((\log N)^{d-1}N^{-1})$. This rate is asymptotically faster than the rate of standard Monte Carlo integration $\mathcal{O}(N^{-1/2})$ for large sample sizes $N$ (Morokoff1994). Nevertheless, importance sampling with space-filling sequences can be inefficient for small-to-medium sample sizes $N$. This is because these sequences are designed for uniform exploration, failing to adapt to the high-probability regions of a specific target. utilizing a space-filling sequence within the BIS framework can be viewed as a sequential permutation of the sequence, such that exploration begins in the high-probability regions of the target. This ensures robust approximation accuracy across both small and large sample sizes $N$.
Our default recommendation of the proposal sequence is the Halton sequence (Halton1964), which is arguably the most widely adopted space-filling sequence in practice. While originally defined for the unit cube, the Halton sequence is readily amenable to arbitrary hyperrectangles. For clarity, we define the Halton sequence scaled to the hyperrectangular domain $\Theta=\prod_{i=1}^{d}[a_{(i)},b_{(i)}]$, where $a_{(i)}$ and $b_{(i)}$ denote the lower and upper boundaries for each coordinate $i$, respectively.
@@
Let $\{\eta_{i}\}_{i=1}^{\infty}$ be the standard Halton sequence on the unit cube $0,1]^{d}$. The scaled Halton sequence $\{\theta_{i}\}_{i=1}^{\infty}$ on the domain $\Theta$ is defined by $\theta_{i}:=a+B\eta_{i}$, where $a:=(a_{(1)},\cdots,a_{(d)})\in\mathbb{R}^{d}$ and $B$ is a $d\times d$ diagonal matrix with $(i,i)$-th entry $b_{(i)}-a_{(i)}$.
Having established the general framework of BIS, we now turn to the analysis of its theoretical guarantees. Subsequently, we detail a practical instantiation of the point-selection criterion.
4 Convergence Guarantee
We present a rigorous analysis of the convergence properties of the BIS framework. Our aim is to highlight that BIS does not compromise the convergence rate of standard importance sampling, regardless of the choice of the point-selection criterion. This resilience stems from the construction, which ensures that no candidate point is selected more than once (c.f. Remark˜1). This intrinsic convergence guarantee grants users the flexibility to design bespoke point-selection criteria, safeguarding against the risk of complex designs introducing unquantifiable sampling biases.
First, Section˜4.1 introduces the Maximum Mean Discrepancy (MMD; Smola2007; Gretton2012), which serves as a metric to assess the approximation error between the weighted samples and the target density. Section˜4.2 then establishes the convergence rate of Algorithm 1 under the default setting of the uniform proposal density. Finally, Section˜4.3 provides the convergence rate of Algorithm 1 in a general setting where the proposal density is arbitrary. Proofs of all the theoretical results are provided in Supplement˜B.
4.1 Metric of Convergence
Consider a metric on the space of probability distributions on $\Theta$ that metrizes the weak convergence; that is, a probability distribution converges weakly to a target if and only if their distance under the metric converges to zero. A standard strategy for establishing the convergence of the weighted samples to the target density is to demonstrate the decay of the metric between them. In this context, the MMD is a computationally and theoretically well-suited choice of the metric.
The MMD is a metric defined via a kernel function $\kappa:\Theta\times\Theta\to\mathbb{R}$. Recall that every kernel implies the existence of the existence of a uniquely associated Hilbert space of functions on $\Theta$, known as the reproducing kernel Hilbert space (RKHS). See, e.g., Paulsen2016 for an introduction to RKHS. Let $\mathcal{H}_{\kappa}$ be the RKHS associated with $\kappa$, equipped with the norm $\|\cdot\|_{\mathcal{H}_{\kappa}}$. The MMD is defined as the worst-case error between two expectations with respect to $p$ and $q$:
The MMD satisfies the metric axioms and metrizes weak convergence of probability distributions under mild regularity conditions on the kernel (Simon-Gabriel2023). For our theoretical analysis, the specific choice of the kernel $\kappa$ is not critical. Provided the condition outlined below, our convergence bounds are independent of the kernel choice. Consequently, the existence of any kernel satisfying this condition suffices to establish the weak convergence.
A kernel $\kappa$ is said to be integrally strictly positive definite if $\int_{\Theta\times\Theta}\kappa(\theta,\theta^{\prime})d\mu(\theta)d\mu(\theta^{\prime})>0$ for all finite non-zero signed Borel measures $\mu$ on $\Theta$. This is analogous to strictly positive definiteness of kernels. For example, Gaussian kernels satisfy this property (Sriperumbudur2010). To streamline our derivation, we impose the following condition on the kernel $\kappa$ used for MMD. Without loss of generality, we assume that both the kernel and the derivatives are bounded by $1$. This is justified because any bounded kernel can be rescaled by a constant factor to satisfy this condition. The Gaussian kernels is a standard example satisfying this condition.
Kernel Condition: The kernel $\kappa:\Theta\times\Theta\to\mathbb{R}$ is integrally strictly positive definite. The kernel $\kappa$ admits a factorization $\kappa(\theta,\theta^{\prime})=\prod_{i=1}^{d}\kappa_{i}(\theta_{(i)},\theta_{(i)}^{\prime})$ for some univariate kernels $\kappa_{1},\dots,\kappa_{d}$, where $\theta_{(i)}$ denotes the $i$-th coordinate of $\theta$. For each $i=1,\dots,d$, the kernel $\kappa_{i}$ and the mixed partial derivative $(\partial/\partial\theta_{(i)})(\partial/\partial\theta^{\prime}_{(i)})\kappa_{i}(\theta_{(i)},\theta^{\prime}_{(i)})$, assumed to exist, are both uniformly bounded by $1$.
4.2 Deterministic Rate in Uniform Case
Our first analysis focuses on the case where the proposal density $u$ is uniform, as discussed in Section˜3.2. Notably, the uniformity of $u$ yields a deterministic convergence rate, irrespective of whether the proposal sequence is stochastic or deterministic. A central quantity in this analysis is the star discrepancy $\operatorname{D}^{*}$ of a set of $K$ points $\{\theta_{n}\}_{n=1}^{K}$ in $\Theta$, defined as
$\operatorname{D}^{*}\left(\{\theta_{n}\}_{n=1}^{K}\right):=\sup_{B\in\mathcal{B}}\left|\frac{\#\{\theta_{n}\in B\}}{K}-\frac{\operatorname{Vol}(B)}{\operatorname{Vol}(\Theta)}\right|,$ (7)
where $\mathcal{B}$ denotes the set of all sub-boxes in $\Theta$ anchored at the lower corner of $\Theta$. Here, $\{\theta_{n}\in B\}$ denotes the number of points contained in the box $B$, and $\operatorname{Vol}(\cdot)$ denotes the volume of a given region. Although the star discrepancy is typically defined for the unit cube, we adopt the above definition extended to the domain $\Theta$.
Let $\partial_{i}f(\theta)$ denote the partial derivative of a function $f(\theta)$ with respect to the i-th coordinate. Let $\partial_{1:d}f(\theta)$ denote the first mixed partial derivative of $f(\theta)$, that is, $\partial_{1:d}f(\theta)=\partial_{1}\cdots\partial_{d}f(\theta)$. Hereafter, let $\delta_{N}$ denote the empirical distribution associated with the $N$ weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ obtained from BIS. We are now in a position to state our main result. The first term in the bound below is derived by invoking the Koksma-Hlawka inequality, a classical result in QMC theory governing numerical integration error. The second term arises from the point-selection procedure of BIS, which decays significantly faster than the canonical Monte Carlo integration rate $\mathcal{O}(N^{-1/2})$.
Theorem 1.
Assume that the target density $p$ admits a uniformly continuous mixed partial derivative $\partial_{1:d}p(\theta)$, and that the proposal density $u$ is uniform. For any proposal sequence $\{\theta_{n}\}_{n=1}^{M+N}$ and selection criterion $U_{n}$, the weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ of Algorithm 1 satisfy
$d(p,\delta_{N})\leq C_{1}\operatorname{D}^{*}\left(\{\theta_{n}\}_{n=1}^{N+M}\right)+C_{2}\frac{M}{N+M},$ (8)
where $C_{1}$ and $C_{2}$ are constants dependent sorely on $p$ and $u$.
Theorem˜1 immediately yields an explicit convergence rate for BIS. To derive this, it suffices to employ a proposal sequence $\{\theta_{n}\}_{n=1}^{M+N}$ with a known upper bound on the star discrepancy. The Halton sequence, for example, satisfies a well-established bound (see e.g. Niederreiter1992). Consequently, we obtain the following rate under the Halton sequence in Definition˜1.
Corollary 1.
Under the assumptions of Theorem˜1, suppose that the proposal sequence $\{\theta_{n}\}_{n=1}^{M+N}$ is the scaled Halton sequence. The weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ satisfy
$d(p,\delta_{N})\leq C_{1}\frac{\log(N+M)^{d}}{N+M}+C_{2}\frac{M}{N+M},$ (9)
where $C_{1}$ and $C_{2}$ are constants dependent solely on $p$ and $u$.
We note that i.i.d. random samples from the uniform density $u$ also satisfy a bound of the star discrepancy that holds almost surely (Morokoff1994). However, we prioritize the deterministic sequence in this work due to its faster convergence rate. Nevertheless, it is worth highlighting that i.i.d. random samples still ensure the almost sure convergence. The established convergence rate implies the consistency of the Monte Carlo estimator by BIS.
Corollary 2.
Under the assumptions in Corollary˜1, the empirical distribution $\delta_{N}$ of the weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ converges weakly to the target density $p$ as $N\to\infty$. Consequently, the resulting Monte Carlo estimator of $\mathbb{E}_{\theta\sim p}[f(\theta)]$ is consistent for any continuous integrand $f$.
4.3 Concentration Rate in General Case
We extend our analysis to a general setting where the proposal density $u$ is arbitrary, and the proposal sequence $\{\theta_{i}\}_{i=1}^{\infty}$ consists of i.i.d. random samples drawn from $u$. Given the stochastic nature of the proposal sequence $\{\theta_{i}\}_{i=1}^{\infty}$, we establish a concentration bound for the MMD.
Theorem 2.
Assume that the proposal density $u$ is uniformly bounded on $\Theta$. Let the proposal sequence $\{\theta_{n}\}_{n=1}^{M+N}$ be i.i.d. samples from $u$. For any selection criterion $U_{n}$, the weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ of Algorithm 1 satisfy that, for any $0<\epsilon<1$,
where $\mathbb{P}$ denotes the probability with respect to the random samples $\theta_{1},\dots,\theta_{M+N}\overset{i.i.d.}{\sim}u$, and $C_{1}$, $C_{2}$, and $C_{3}$ denote constants dependent sorely on $p$ and $u$.
This theorem leads to an immediate corollary regarding the almost-sure weak convergence of the weighted samples, and hence the consistency of the Monte Carlo estimator by BIS.
Corollary 3.
Under the assumptions in Theorem˜2, the empirical distribution $\delta_{N}$ of the weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ converges weakly to the target density $p$ almost surely as $N\to\infty$. Consequently, the resulting Monte Carlo estimator of $\mathbb{E}_{\theta\sim p}[f(\theta)]$ is consistent for any continuous integrand $f$.
Although this general setting accommodates arbitrary proposal densities, the resulting convergence rate is stochastic. Furthermore, the deterministic rate established in Corollary˜1 is notably faster in lower dimensions $d$. This observation underscores the advantage of using the Halton sequence in Definition˜1 as the proposal sequence.
5 Adapting Bayesian optimization for Sampling
We now present a versatile, practical instance of the point-selection criteria, which renders BIS effective across a range of black-box densities in low dimensions. This criterion leverages GP-surrogate modeling, analogous to BO. In machine learning, BO has established itself as a powerful approach to the optimization of black-box functions. BIS enables the application of this powerful bandit approach to sampling.
As mentioned in Section˜4, BIS allows for great flexibility in designing bespoke point-selection criteria. The development of further variants beyond BO is reserved for future work, as the rigorous derivation of each new criterion necessitates a dedicated study, and providing an exhaustive catalog of point-selection criteria lies beyond the scope of this paper.
@@ 5.1 Point-Selection Criterion
The point-selection criterion $U_{n}$ focused in this paper leverages a GP surrogate model for the unnormalized density $q(\theta)$, the proportional term of the target density $p(\theta)$. The unnormalized density $q(\theta)$ is strictly positive, while GP functions take values in the real line $\mathbb{R}$. To fit a GP to the density, a link function to map the density to the real line is required. The log-transformation, $\log q(\theta)$, is a standard approach. Here, we generalize this by employing a convex non-negative function $\phi:\mathbb{R}\to[0,\infty)$ whose inverse $\phi^{-1}$ is well-defined on $(0,\infty)$.
Figure 2: Illustration of GP-UJB using the unnormalized standard normal target $q$ and the exponential map $\phi(\cdot)=\exp(\cdot)$. Left: The transformed GP posterior $\phi(f(\theta))$ conditional on five evaluations of $q$. Center: The resulting GP-UJB, the selected location for the next evaluation, and the decomposition into the exploitation and exploration terms. The exploitation term drives selection towards the mode of the current estimate, while the exploration term drives selection towards where $\phi(f(\theta))$ is most uncertain. Right: The transformed GP posterior after the new evaluation.
Under this formulation, we model the transformed unnormalized density $g(\theta)=\phi^{-1}(q(\theta))$ via a GP, where the fitted GP can be transformed back to the non-negative range by $\phi(\cdot)$. For example, the log-transformation $\phi^{-1}(\cdot)=\log(\cdot)$ corresponds to the choice $\phi(\cdot)=\exp(\cdot)$. Now, we introduce our criterion, GP-UJB, equipped with appealing properties for multi-armed bandits.
@@
$U_{n}(\theta):=\mathbb{E}_{f\sim\mathcal{GP}(m_{n},k_{n})}\left[\phi\big(f(\theta)\big)\right],$ (11)
where $f(\theta)$ denotes the GP sample path $f\sim\mathcal{GP}(m_{n},k_{n})$ evaluated at $\theta$.
Driving both exploitation of past information and exploration of unexplored domains is a pivotal factor in most bandit problems. Notably, GP-UJB admits an interpretable decomposition into two terms that drive, respectively, exploitation and exploration.
Remark 2 (Exploitation-Exploration Decomposition).
GP-UJB can be decomposed as
@@
The first term is the best predictor based on current information, i.e., the least-error plug-in approximation of $q(\theta)$ attainable by the GP. The second term is known as the Jensen’s gap and non-negative due to Jensen’s inequality. It measures the amount of uncertainty in $\phi(f(\theta))$ at each location $\theta$. This term is zero at the training points $\{\theta_{i}\}_{i=1}^{n}$, where the GP posterior has no uncertainty. Figure˜2 illustrates the decomposition of GP-UJB computed for a standard normal density in $\mathbb{R}$.
5.2 Approximation Error of GP-UJB
We establish that the L2 error between GP-UJB and the unnormalized target $q(\theta)$ is controlled by the average-case interpolation error of the GP regression. In particular, if an additional regularity condition is met, the former is specifically smaller than the later. This also supports the use of GP-UJB as it converges faster than the concentration of the GP regression itself. In the following result, $\|\cdot\|_{L_{2}(\Theta)}$ denotes the L2 norm of functions on $\Theta$ under the Lebesgue measure.